Find the solution of the differential equation that satisfies the given initial condition. dL = KL2 In(t), L(1) = -1 dt X
Answer #1

· 거+= 3 + = F <- D+[9] = F + [F - (FIUTFTA = 17- G F-17 - + [F - (1)Uta = 14 G t ty 아 dy 가가 것 ot) 해신 klen(t) de 기 F L k en LSI = el text)-t]+ k +1.W = kt-ktinct) - k-1 k is LE kt-ktent) -K